Y = 1/ (a-z)^2, find dy/dz My answer is [-2(a-z)^(-3) ] Can anyone please rectify this and elaborate

Grade:10

1 Answers

Vikas TU
6 years ago
We know that d/dx(1/x2)=-2/x3
Similaryly,
d/dz(1/(a-z)^2)= -2(a-z)^(-3).
Keeping the a as a constant.